OpenStudy (anonymous):

\[\lim_{x \rightarrow + \infty} ( \frac{ x }{ x-1 } - \frac{ 1 }{ \ln (x) })\]

OpenStudy (shubhamsrg):

at x-> inf, 1/ln(x) is ofcorse 0 so your main problem must be with x/(x-1) right?

OpenStudy (anonymous):

yes

Parth (parthkohli):

L'Hopital baby!

OpenStudy (shubhamsrg):

You must be knowing LH rule ?

OpenStudy (shubhamsrg):

lol..

OpenStudy (anonymous):

yes i know

OpenStudy (shubhamsrg):

So, where are you stuck then ?

Parth (parthkohli):

\[\dfrac{d}{dx} x = \cdots\]and\[\dfrac{d}{dx} x - 1 = \cdots\]Though I first realized the answer through a good trick.

OpenStudy (anonymous):

i want answer only

OpenStudy (shubhamsrg):

If you want the answer,then get your answer! ;)

Parth (parthkohli):

Basically, if you look at the fraction \(\dfrac{x}{x -1 }\), you will realize that when \(x\) grows, then \(\dfrac{x}{x - 1}\) grows too. But it never really reaches \(1\), but it does keep approaching it!

OpenStudy (shubhamsrg):

you may even treat x/(x-1) as (x-1)+1 /(x-1) = 1 + 1/(x-1) now put x->inf.

Parth (parthkohli):

Like \(x = 2 \iff \dfrac{x}{x - 1} = 2\). \[x = 1000 \iff \dfrac{x}{x - 1} = \dfrac{1000}{1001} \]It's getting closer and closer to \(1\).

OpenStudy (shubhamsrg):

Well thats the whole importance of limits @ParthKohli rightly says here. Suppose we have x-2 = x + 9 and we want to solve for x, Many will simply say no solution; but if you talk about limits, x=inf is one answer. Limits is all theoritical. Because of limits only we can say that 2 parallel lines meet at infinity or vice versa.

OpenStudy (shubhamsrg):

Your case is similiar x and x-1 are parallel lines. but x=x-1at x->inf hence the ratio =1 at inf

Parth (parthkohli):

Yes, I like theoretical stuff 8-)